# The Value of Information in Human-AI Decision-making

### v5

@@ E.2 Results
We first analyze the agent-complementary information values in Figure 4, using the Brier score as the payoff function. We find that all AI models provide complementary information value to the aforementioned human judgment. As shown in Figure 4 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions. This motivates deploying an AI to assist humans in this scenario.
In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 4. This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
@@
Figure 4 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure 6. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
Appendix F Demonstration II: Behavioral Analysis on Deepfake Detection
@@
We apply our framework to analyze a deepfake video detection task [Dolhansky et al., 2020], where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
@@
Figure 5: Information value calculated under our framework in the information model defined by the experiment of Groh et al. [2022]. Basic signals include the seven video level features and three types of agent decisions. The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
F.1 Data and Model
@@
We define the information model on the experiment data of Groh et al. [2022]. Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model. Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network [Zhang et al., 2016], with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision. This produced a final decision ($D^{\text{Human-AI}}$). Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
@@
We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from the mean absolute error used by Groh et al. [2022], but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
@@
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. [2022]: graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Appendix E while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
F.2 Results
@@
We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance). Figure 5(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$). This is consistent with the findings of Guo et al. [2024] that humans are bad at distinguishing when AI predictions are correct.
@@
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions. As shown on the fifth row in Figure 5, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure 5, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
@@
By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions. As shown in Figure 5, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face. This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
Appendix G Robustness Analysis in Demonstration I
@@
Appendix H Screenshots of the Experiment

### v6

@@ F.2 Results
We first analyze the agent-complementary information values in Figure˜5, using the Brier score as the payoff function. We find that all AI models provide complementary information value to the aforementioned human judgment. As shown in Figure˜5 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions. This motivates deploying an AI to assist humans in this scenario.
In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure˜5. This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
@@
Figure˜5 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition˜B.3, we test the payoff of models on all V-shaped scoring rules, shown in Figure˜7. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
Appendix G Demonstration II: Behavioral Analysis on Deepfake Detection
@@
We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
@@
Figure 6: Information value calculated under our framework in the information model defined by the experiment of Groh et al. (2022). Basic signals include the seven video level features and three types of agent decisions. The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
G.1 Data and Model
@@
We define the information model on the experiment data of Groh et al. (2022). Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model. Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision. This produced a final decision ($D^{\text{Human-AI}}$). Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
@@
We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from the mean absolute error used by Groh et al. (2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
@@
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Appendix˜F while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.
G.2 Results
@@
We first compare the information value of the AI predictions to that of the human decisions in the first round (without AI assistance). Figure˜6(a) shows that AI predictions provide about $65\%$ of the total possible information value over the no-information baseline, while human decisions only provide about $15\%$. Because the no information baseline, $0.75$, is equivalent to a random guess drawn from $\text{Bernoulli}(0.5)$, human decisions are only weakly informative for the problem.
We next consider the human-AI decisions. Given that the AI predictions contain a significant portion of the total possible information value, we might hope that when participants have access to the AI predictions, their performance will be close to the full information baseline. However, the information value of the human-AI decisions only achieves a small proportion of the total possible information value ($30\%$). This is consistent with the findings of Guo et al. (2024) that humans are bad at distinguishing when AI predictions are correct.
@@
To understand what information might improve human decisions, we assess the $\operatorname{ACIV}$s of different signals over different agents. This describes the additional information value in the signal after conditioning on the existing information in the agents’ decisions. As shown on the fifth row in Figure˜6, the presence of a flickering face offers larger $\operatorname{ACIV}$ over human decisions than over AI predictions, meaning that human decisions could improve by a greater amount if they were to incorporate this information. Meanwhile, as shown on the fourth row in Figure˜6, the presence of an individual with dark skin offers larger $\operatorname{ACIV}$ over AI predictions than over human decisions, suggesting that humans make greater use of this information. This suggests that the AI and human rely on differing information to make their initial predictions, where the AI relies more on information associated with the presence of a flickering face while human participants rely more on information associated with the presence of an individual with dark skin.
@@
By comparing the $\operatorname{ACIV}$s of different signals over human decisions and human-AI decisions, we also find that simply displaying AI predictions to humans did not lead to the AI-assisted humans exploiting the observed signals in their decisions. As shown in Figure˜6, with the assistance of AI, the $\operatorname{ACIV}$s of all signals over the human-AI teams’ decisions do not change much compared to the $\operatorname{ACIV}$s over human decisions, with the exception of a slight improvement in the presence of a flickering face. This finding further confirms the hypothesis that humans are simply relying on AI predictions without processing the information contained in them.
Appendix H Sensitivity Analysis of the Rational Belief Estimator
In this section, we first present a theoretical analysis on the quality of the ACIV estimator by connecting the calibration error of the given $\mathcal{A}$. Then, we present an empirical analysis on the chest X-ray diagnosis demonstration with different algorithms (linear regression (LR), gradient boosting methods (GBM), and neural network (NN)) as $\mathcal{A}$.
H.1 Theoretical Analysis
In this section, we theoretically show how the quality of the ACIV estimator (Algorithm˜1) is related to the calibration error of the rational belief estimator $a=\mathcal{A}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$. We use the regret of the decision maker to represent the quality of the estimation on the rational decision rules, i.e., how much payoff can be improved if we correct the decisions in hindsight [^12] We use the notion of regret to quantify the quality of our ACIV estimator because the perfect decisions are usually unidentifiable with finite signals.. Because the ACIV is defined as the improvement of the best-attainable performance with the signal and the agent decision over the one with the agent decision alone, the decisions $\hat{d}^{r}$ and $\hat{d}^{rb}$ in Algorithm˜1 should have as small regret as possible. We first derive our definition of the regret of Algorithm˜1 using the swap regret by Roth and Shi (2024).
Definition H.1 (Swap Regret (Roth and Shi, 2024)).
Given a set of observations $\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n}$, a decision rule $d(\cdot)$, and a decision task with payoff function $S$, the swap regret of the DM is:
$\textsc{Swap}_{S}(d,\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})=\frac{1}{n}\max_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(d(v_{i},d^{b}_{i})),\omega_{i})-S(d(v_{i},d^{b}_{i}),\omega_{i})\right]$ (11)
[…]
Definition H.2 (Regret of the ACIV estimator).
Given an empirical distribution as the data-generating process, $\pi=\text{Uniform}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$, the estimated decision rules $\hat{d}^{r}:\mathbf{V}\times\mathbf{D}\rightarrow\mathbf{D}$ and $\hat{d}^{rb}:\mathbf{D}\rightarrow\mathbf{D}$ from Algorithm˜1, and a decision task with payoff function $S$, the regret of the ACIV estimator is:
$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})=\operatorname{ACIV}^{\pi,S,d^{*r},d^{*rb}}(V;D^{b})-\operatorname{ACIV}^{\pi,S,\hat{d}^{r},\hat{d}^{rb}}(V;D^{b})$ (12)
where $\operatorname{ACIV}^{\pi,S,\hat{d}^{r},\hat{d}^{rb}}$ denotes the $\operatorname{ACIV}$ estimated under the rational decision rules $\hat{d}^{r}$ and $\hat{d}^{rb}$. $d^{*r}$ and $d^{*rb}$ are the optimal decision rules that we can get from $\hat{d}^{r}$ and $\hat{d}^{rb}$:
$d^{*r}(\cdot,\cdot)=\sigma^{r}(\hat{d}^{r}(\cdot,\cdot))\text{ where }$ $\sigma^{r}=\operatorname*{arg\,max}_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{r}(v_{i},d^{b}_{i})),\omega_{i})\right]$
$d^{*rb}(\cdot)=\sigma^{rb}(\hat{d}^{rb}(\cdot))\text{ where }$ $\sigma^{rb}=\operatorname*{arg\,max}_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{rb}(d^{b}_{i})),\omega_{i})\right]$
Lemma H.3.
The regret of the ACIV estimator is upper bounded by the swap regret of $\hat{d}^{r}$:
$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})\leq\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (13)
Proof.
$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})$ $=\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\sigma^{r}(\hat{d}^{r}(v,d^{b})),\omega_{i})}]-\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(v,d^{b}),\omega_{i})}]$- (14)
$\quad\left(\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(\sigma^{rb}(\hat{d}^{r}(d^{b}_{i})),\omega_{i})}]-\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(d^{b}_{i}),\omega_{i})}]\right)$
$\leq\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\sigma^{r}(\hat{d}^{r}(v,d^{b})),\omega_{i})}]-\text{\bf E}_{(v,d^{b},\omega)\sim\pi}\![{S(\hat{d}^{r}(v,d^{b}),\omega_{i})}]$
$=\max_{\sigma:\mathbf{D}\rightarrow\mathbf{D}}\left[\frac{1}{n}\sum_{i=1}^{n}\left[S(\sigma(\hat{d}^{r}(v_{i},d^{b}_{i})),\omega_{i})-S(\hat{d}^{r}(v_{i},d^{b}_{i}),\omega_{i})\right]\right]$
$=\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$
Claim H.4 (Kleinberg et al. (2023), Theorem 12).
Given a decision task with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\rightarrow[0,1]$, if the DM responds by taking $d(v,d^{b})=d^{*}(a(v,d^{b}))$, where $a$ is an estimator of the probability of the payoff state given the signal and action and $d^{*}(p)=\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$ is the best response to the probability $p$, the swap regret of the DM is bounded by the expected calibration error (ECE) of the estimator $a$:
$\textsc{Swap}_{S}(\hat{d}^{r},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})\leq 2\text{ECE}(a,\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (15)
Theorem H.5.
Given a decision task with bounded payoff $S:\mathbf{D}\times\mathbf{\Omega}\rightarrow[M_{1},M_{2}]$, the regret of the ACIV estimator is bounded by the expected calibration error of the estimator $\hat{a}=\mathcal{A}(\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n})$:
$\textsc{RegACIV}_{S,\pi}(\hat{d}^{r},\hat{d}^{rb};V,D^{b})\leq 2(M_{2}-M_{1})\textsc{ECE}(\hat{a},\{(v_{i},d^{b}_{i},\omega_{i})\}_{i=1}^{n})$ (16)
The proof normalizes the payoff in Lemma˜H.3 into $[0,1]$ and then applies ˜H.4.
Theorem˜H.5 shows that, in Algorithm˜1, when we choose a predictive algorithm $\mathcal{A}$ that yields low ECE, the ACIV estimated by Algorithm˜1 is close to the ACIV under the optimal decision rules.
H.2 Empirical Analysis
In this section, we take the chest X-ray diagnosis task in Appendix˜F as an example to empirically analyze the estimation of the full information value by Algorithm˜1 with different modeling approaches (linear regression (LR), gradient boosting methods (GBM), and neural network (NN)).
Task and data.
We use the chest X-ray diagnosis task in Appendix˜F as the example. The task is to predict the probability of the presence of a disease given a chest X-ray image. The decision space is $\mathbf{D}=\Delta[0,1]$ and the payoff space is $\mathbf{\Omega}=\{0,1\}$. The payoff function is $S(d,\omega)=1-(d-\omega)^{2}$. The signal space is a high-dimensional feature vector of the chest X-ray image. We split the dataset that we used in Appendix˜F into a 70/30 train/test split.
Predictive Models.
We use the following models:
  • Linear Regression (LR) from scikit-learn package.
  • Extreme Gradient Boosting (GBM) from xgboost package with n_estimators = 100, max_depth = 6, and learning_rate = 0.3.
  • MLP classifier (NN) from scikit-learn package with 2 hidden layers of 256 and 64 neurons each.
Evaluation Metrics.
We report the following evaluation metrics for each estimator:
  • Brier Score: $\frac{1}{n}\sum_{i=1}^{n}(\hat{p}_{i}-\omega_{i})^{2}$
  • Accuracy: $\frac{1}{n}\sum_{i=1}^{n}\mathbb{1}(\mathbb{1}(\hat{p}_{i}\geq 0.5)=\omega_{i})$
  • Expected Calibration Error (ECE): $\frac{1}{n}\sum_{i=1}^{n}\left|\hat{p}_{i}-\text{\bf E}\![{\omega|\hat{p}_{i}}]\right|$
Table 7: Performance of different algorithms for ACIV estimation in the chest X-ray diagnosis task. Metrics for the estimators include Brier Score, Accuracy, F1-Score, and Expected Calibration Error (ECE). We report the estimated $\operatorname{ACIV}(V;\emptyset)$ for each estimator with Algorithm˜1.
Model Brier Score $\uparrow$ Accuracy $\uparrow$ ECE $\downarrow$ $\mathbf{ACIV(V;\emptyset)}\uparrow$
Linear Regression (LR) 0.834 0.775 0.039 0.086
Gradient Boosting (GBM) 0.841 0.771 0.026 0.108
Neural Network (NN) 0.853 0.789 0.036 0.0757
Results.
The results are shown in Table˜7. We can see that the GBM estimator achieves the best performance in terms of ECE, while the NN estimator achieves the best performance in terms of Brier Score and Accuracy. This validates our theoretical result: ECE is a good proxy for the regret of the estimation of the ACIV by Algorithm˜1 compared to the Brier Score and Accuracy.
Appendix I Robustness Analysis in Demonstration I
@@
Appendix J Screenshots of the Experiment